How many liters of distilled water must be added to 140 liters of an 80% alcohol solution to obtain a 70% solution? i think the answer is 2.3 liters.
Answer by mananth(16946)

------ Percent ---------------- quantity
Alcohol 80 ---------------- 140
water 0 ---------------- x
Total 70 ---------------- 140+x
80*140+ 0*x=70(140+x)
11200=9800+70x

-70x=9800-11200
-70x=-1400
/-70x
x=20 gallons water